OpenStudy (anonymous):

8. Consider the vector v = <3, 2>. a. What is the angle between v = <3, 2> and the positive x-axis? b. Write v in the form (|v|cosθ|v|sinθ).

OpenStudy (amistre64):

dot product may help

OpenStudy (amistre64):

\[\cos\alpha=\frac{\vec u\cdot \vec v}{|\vec u|~|\vec v|}\]

OpenStudy (anonymous):

can you explain it? i will be back have to eat lunch

OpenStudy (amistre64):

let u = <1,0> for simplicities sake: |u| = 1, and u.v = xpart of v

OpenStudy (amistre64):

i play in real time .... so when you get back we will discuss it if im available

OpenStudy (anonymous):

@amistre64